# Efficient Diffusion Models: A Survey

### v1

@@ 3.1 Efficient Training
Efficient training techniques focus on reducing the costs of the DM pre-training process in terms of compute resources, training time, memory and energy consumption. As summarized in Figure 2, enhancing the efficiency of pre-training can be achieved through different and complementary techniques, including noise schedule, score matching, data-dependent adaptive priors, and rectified flow.
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=18em,font=,, where level=2text width=15em,font=,, where level=3text width=24em,font=,, [ Efficient Training, ver [ Noise Schedule, fill=blue!10 [ Fixed Noise Schedule, fill=blue!10 [ DDPM (Ho et al., 2020), IDDPM (Nichol & Dhariwal, 2021), Laplace (Hang & Gu, 2024),
Masked-Diffusion LM (Chen et al., 2023a), leaf, text width=45em ] ] [ Adaptive Noise Schedule, fill=blue!10 [ DDIM (Song et al., 2020a), ResShift (Yue et al., 2024), Immiscible Diffusion (Li et al., 2024d),
DiffuSeq (Gong et al., 2022), SeqDiffuSeq (Yuan et al., 2024), leaf, text width=45em ] ] ] [ Score Matching, fill=blue!10 [ Score Matching (Hyvärinen & Dayan, 2005), Sliced Score Matching (Song et al., 2020b), NCSN (Song & Ermon, 2019),
Likelihood Weighting (Song et al., 2021), CLD (Dockhorn et al., 2021), leaf, text width=61.7em ] ] [ Data-Dependent Adaptive Priors, fill=blue!10 [ PriorGrad (Lee et al., 2021), Digress (Vignac et al., 2022), DecompDiff (Guan et al., 2023), Leapfrog DM (Mao et al., 2023), leaf, text width=61.7em ] ] [ Rectified Flow, fill=blue!10 [ Rectified Flow (Liu et al., 2022b; Liu, 2022), InstaFlow (Liu et al., 2023b), PeRFlow (Yan et al., 2024),
2-Rectified Flows++ (Lee et al., 2024a), SlimFlow (Zhu et al., 2024), leaf, text width=61.7em ] ] ]
@@
3.1.1 Noise Schedule
Noise schedule is a crucial component of diffusion models, governing how noise is added at each step of the forward process and how it is removed during the reverse process. Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020) first proposed a linear noise schedule that gradually decreases the variance of the noise added in the forward process in  Eq.(1). However, despite DDPM demonstrating significant potential in diffusion model tasks, this linear schedule requires calculating complex noise terms and iterating through numerous timesteps, posing challenges to noise injection efficiency and highlighting the need for an efficiency-enhanced noise schedule design. Current studies still face prolonged iterative processes, significantly hindering diffusion speed, while several studies have concentrated on designing the highlight of noise schedules. As shown in Figure 3, efficient noise schedules can be classified into fixed noise schedule and adaptive noise schedule.
Figure 3: Illustration of two categories of noise schedules.
Fixed Noise Schedule. Fixed noise schedules refer to a technique where noise is systematically added to a model’s training process at predefined intervals or according to specific levels. DDPM (Ho et al., 2020) uses a linear noise schedule, serving as a classic example of fixed noise schedules, with noise variance changing deterministically over time. Based on this, the Improved Denoising Diffusion Probabilistic Model (IDDPM) (Nichol & Dhariwal, 2021) introduces significant innovations in the noise schedule, replacing the linear noise schedule of the original DDPM with a cosine noise schedule, which could be written as
$\beta_{t}=1-\frac{\cos\left(\frac{t/T+s}{1+s}\cdot\frac{\pi}{2}\right)}{\cos\left(\frac{s}{1+s}\cdot\frac{\pi}{2}\right)}$ (7)
where $t$ is the current timestep, $T$ is the total number of timesteps, and $s$ is a small positive constant typically used to smooth the initial noise addition. The cosine noise schedule helps the model achieve high-quality sample generation by optimizing the noise distribution throughout the training process, allowing the model to generate samples with a reduced number of sampling steps while preserving essential structural information at each stage. However, the cosine noise schedule allocates computational resources evenly across the entire range of noise intensities. This means that both high and low noise samples consume significant resources, even though they may not be the most critical for training the model.
To address the inefficiency in resource allocation,  (Hang & Gu, 2024) proposed a fixed noise schedule, called Laplace. This schedule aims to further optimize the training of the diffusion model by increasing the sampling frequency around critical regions. This approach ensures that the model receives more training samples in these crucial regions, thereby improving overall training efficiency and sample quality. By redesigning the fixed noise schedule, more computational resources are concentrated on medium noise intensities. These medium noise samples are more important for model learning, as they effectively train the model to understand the data structure and remove noise. Laplace effectively balances noise addition across all time steps, ensuring a more robust and consistent training process, and ultimately resulting in higher-quality generated images.
Lastly, existing diffusion models have employed a uniform addition of Gaussian noise to each word to facilitate the diffusion process for text generation. However, this approach has been shown ta failure to leverage the linguistic features of the text in question, while simultaneously increasing the computational burden. Furthermore,  (Chen et al., 2023a) addresses this issue by gradually introducing noise through a soft-masking noise strategy. This model determines the masking order based on word importance, which is measured by term frequency and information entropy. The model employs a square-root noise schedule(Li et al., 2022) to incrementally increase the noise level, thereby stabilizing the training process. Consequently, the model gradually adds noise to the initial latent variable $X_{0}$, resulting in a series of noisy latent variables $X_{1:T}$.
Adaptive Noise Schedule. Different from the fixed noise schedules, adaptive noise schedules include methods that dynamically adjust the noise schedule based on the state of the model or data. The key idea is to adapt the noise schedule in response to specific conditions or states during the diffusion process.  Song et al. (2020a) developed Denoising Diffusion Implicit Models (DDIM), which improve the noise schedule in DDPM by introducing non-Markovian forward processes. as follows:
$x_{t-1}=\sqrt{\alpha_{t-1}}\left(x_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}(x_{t})\right)+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\epsilon_{\theta}(x_{t})+\sigma_{t}\epsilon_{t}$ (8)
where $\alpha_{t}$ is a decreasing schedule controlling how noise is added over time. They propose a dynamic method for adjusting the noise level, $\sigma_{t}$, at each step. The noise level is calculated as a function of both the current state, $x_{t}$, and the initial state, $x_{0}$, leveraging the entire trajectory rather than treating each step independently. This dynamic adjustment fully utilizes the information from the initial and current states, allowing for more accurate control of the diffusion process. As a result, the quality of the generated samples is enhanced, and the number of sampling steps required is reduced.
Inspired by DDIM,  Yue et al. (2024) proposed ResShift, introducing a novel noise schedule that constructs a much shorter Markov chain. The key idea is to shift the residual between the high-resolution (HR) and low-resolution (LR) images instead of adding Gaussian noise. This approach allows the model to start from an LR image and iteratively refine it to produce the HR image. The noise schedule also involves a hyperparameter $\kappa$, which controls the overall noise intensity during the transition. The mathematical formulation of this noise schedule is as follows:
$\sqrt{\eta_{t}}=\sqrt{\eta_{1}}\times b_{0}^{\beta_{t}},\quad t=2,\ldots,T-1$ (9)
where
$\beta_{t}=\left(\frac{t-1}{T-1}\right)^{p}\times(T-1),\quad b_{0}=\exp\left(\frac{1}{2(T-1)}\log\frac{\eta_{T}}{\eta_{1}}\right)$ (10)
where $T$ is the total number of timesteps, $t$ is the current timestep, $p$ is a hyperparameter controlling the growth rate of $\sqrt{\eta_{t}}$, and $\eta_{1}$ and $\eta_{T}$ represent the initial and final noise levels, respectively. This formula enables a non-uniform geometric progression of noise levels. Thus ResShift can achieve high-quality SR results with only 15 sampling steps. Conventional methodologies tend to diffuse each image across the entirety of the noise space, thereby resulting in a composite of all images at each point within the noise layer. In light of the aforementioned,  Li et al. (2024d) proposed Immiscible Diffusion, a novel approach inspired by the physical phenomenon of immiscibility. In contrast to conventional methodologies, Immiscible Diffusion involves reassigning noise to images in a way that minimizes the distance between image-noise pairs within a mini-batch, ensuring that each image is only diffused to nearby noise points. This method ensures that each image is matched only with nearby noise, thereby reducing the difficulty of denoising. Previous studies  Gong et al. (2022) on text generation have employed fixed noise schedules and encoder-only Transformer architectures. This approach necessitated the recalculation of the input sequence at each time step, resulting in the inefficient utilization of computational resources and the generation of outputs at a slow pace. In contrast,  Yuan et al. (2024) introduces an adaptive noise scheduling technique that enables the dynamic adjustment of the noise level at each time step and token position. In particular, this technique entails recording loss values at each timestep throughout the training phase, with a linear interpolation subsequently employed to map these losses to the corresponding noise schedule parameters.
3.1.2 Score Matching
While DDPM directly optimizes the noise prediction in Eq.(6), score matching objectives can directly be estimated on a dataset and optimized with stochastic gradient descent, the loss function for score matching takes a different approach formulated as follows:
@@ 3.1.2 Score Matching
Compared to DDPM’s straightforward optimization in Eq.(6), although score matching Eq.(13) avoids the computation of the partition function $\bm{Z}_{\bm{\theta}}$, it still faces computational challenges, particularly in high-dimensional data. The computation of the trace of the Hessian matrix substantially increases the complexity as the dimensionality grows. Specifically, computing the trace requires many more backward passes than the gradient, making score matching computationally expensive for high-dimensional data.
@@
To address this issue, Song et al. (2020b) observed that one-dimensional problems are typically much easier to solve than high-dimensional ones. Inspired by the idea of the Sliced Wasserstein Distance (Rabin et al., 2012), they proposed Sliced Score Matching. The core idea of sliced score matching is to project both the score function of the model $\bm{s}_{m}(\mathbf{x};\bm{\theta})$ and the data $\bm{s}_{d}({\mathbf{x}})$ onto a random direction $\mathbf{v}$, and compare the differences along that direction. The sliced score matching objective is defined as:
@@
Figure 4: Illustration of data-dependent adaptive priors for diffusion processes across different modalities, demonstrating how tailored priors improve generate on quality.
@@
While effective for dimensionality reduction, score estimation still faces challenges in low data density regions where data samples are sparse. Building upon sliced score matching, to address the issue of inaccurate score estimation in low data density regions, Song & Ermon (2019) introduces a novel generative framework that employs Langevin dynamics to produce samples based on estimated gradients of the data distribution $p_{\text{data}}(\mathbf{x})$. They proposed Noise Conditional Score Networks (NCSN) $s_{\theta}(\mathbf{x},\sigma)$, which jointly estimate scores across multiple noise-perturbed data distributions. By conditioning on a geometric sequence of noise levels $\sigma_{3}>\sigma_{2}>\sigma_{1}$, a single network learns to estimate scores for distributions ranging from highly smoothed $p_{\sigma_{3}}(\mathbf{x})$ that fill low-density regions to concentrated $p_{\sigma_{1}}(\mathbf{x})$ that preserve the structure of the original data manifold. This unified training approach enables robust score estimation across the entire data space.
Different from sliced score matching, Song et al. (2021) introduces a new weighting scheme called likelihood weighting, which allows the weighted combination of score matching losses to actually bound the negative log-likelihood of the diffusion models. Compared with traditional score matching, it focuses on adjusting the training objective. This modification enables approximate maximum likelihood training, rather than merely minimizing the score matching loss, which improves the model’s likelihood across various datasets, stochastic processes, and architectures, which effectively combines the training efficiency of score matching with the advantages of maximum likelihood estimation. Following a similar derivation, as Song et al. (2021), Dockhorn et al. (2021) introduces Coupled Langevin Dynamics (CLD), redefining the score matching objective within the CLD framework. Unlike traditional score matching methods that inject noise directly into the data space, CLD simplifies the task by only requiring the model to learn the score of the conditional distribution $p_{t}(v_{t}\mid x_{t})$, where noise is injected into an auxiliary variable $v_{t}$ coupled with the data.
3.1.3 Data-Dependent Adaptive Priors

### v2

@@ 3.1 Efficient Training
Efficient training aims to optimize the training process of DMs, reducing computational costs while accelerating convergence. As summarized in Figure 3, enhancing the efficiency of pre-training can be achieved through different and complementary techniques, including latent diffusion, loss formulation, and specialized training tricks. Latent diffusion models compress the optimization process by operating in lower-dimensional latent spaces, though they occasionally struggle with fine-grained detail reconstruction. Loss formulation methods enhance gradient estimation and stability, though sometimes facing numerical challenges. Meanwhile, various training tricks, including data-dependent adaptive priors, extra loss supervision, and optimized noise schedules, further enhance efficiency by leveraging problem-specific knowledge, though each introduces additional hyperparameters requiring careful tuning.
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=9em,font=,, where level=2text width=17em,font=,, where level=3text width=24em,font=,, [ Efficient Training, ver [ Latent Diffusion, fill=blue!8 [ Video LDM (Blattmann et al., 2023), MagicVideo (Zhou et al., 2022), PVDM (Yu et al., 2023), LVDM (He et al., 2022),
Latent 3D Graph Diffusion (You et al., 2024), Motion Latent-based Diffusion (Chen et al., 2023c), AudioLDM (Liu et al., 2023a),
L3DG (Roessle et al., 2024), TabSyn (Zhang et al., 2023a), PolicyDiffusion (Hegde et al., 2023), RNAdiffusion (Huang et al., 2024), leaf, text width=61.3em ] ] [ Loss Formulation, fill=blue!8 [ Score Matching, fill=blue!10 [ Score Matching (Hyvärinen & Dayan, 2005), Sliced Score Matching (Song et al., 2020b),
NCSN (Song & Ermon, 2019), Likelihood Weighting (Song et al., 2021),
CLD (Dockhorn et al., 2021), leaf, text width=42.7em ] ] [ Rectified Flow, fill=blue!10 [ Rectified Flow (Liu et al., 2022b; Liu, 2022), InstaFlow (Liu et al., 2023c),
PeRFlow (Yan et al., 2024), 2-Rectified Flows++ (Lee et al., 2024a), SlimFlow (Zhu et al., 2024), leaf, text width=42.7em ] ] ] [ Training Tricks, fill=blue!10 [ Data-Dependent Adaptive Priors, fill=blue!10 [ PriorGrad (Lee et al., 2021), Digress (Vignac et al., 2022), DecompDiff (Guan et al., 2023),
Leapfrog DM (Mao et al., 2023), leaf, text width=42.7em ] ] [ Extra Loss Supervision, fill=blue!10 [ REPA (Yu et al., 2024a), VA-VAE (Yao et al., 2025), leaf, text width=42.7em ] ] [ Noise Schedule, fill=blue!10 [ DDPM (Ho et al., 2020), IDDPM (Nichol & Dhariwal, 2021), Laplace (Hang & Gu, 2024),
Masked-Diffusion LM (Chen et al., 2023a), DDIM (Song et al., 2020a),
ResShift (Yue et al., 2024), Immiscible Diffusion (Li et al., 2024d),
DiffuSeq (Gong et al., 2022), SeqDiffuSeq (Yuan et al., 2024), leaf, text width=42.7em ] ] ] ] ]
@@
3.1.1 Latent Space
Early diffusion models, such as DDPM  (Ho et al., 2020), operate directly in pixel space, where images are generated through iterative noise addition and denoising. While these models achieve high-quality results, their reliance on pixel space introduces significant inefficiencies. Each diffusion step involves operations in a high-dimensional space, leading to substantial computational and memory overhead, especially for high-resolution images (e.g., 512×512). Additionally, the requirement for hundreds to thousands of iterative steps hinders real-time applications. Although methods like DDIM  (Song et al., 2020a) have been proposed to speed up sampling, the pixel-level processing remains a fundamental bottleneck. In contrast, leveraging latent space significantly enhances training efficiency by operating in a lower-dimensional, compact representation, reducing both computational complexity and memory consumption while maintaining high-quality generation.
As research progresses, researchers have begun exploring compromise approaches to circumvent the high-dimensional pixel space. Consequently, various methods for introducing the latent space have increasingly gained attention from the academic community. To reduce computational complexity, researchers have introduced autoencoders (AEs) and their variants to compress images into a lower-dimensional latent space. Among them, Variational Autoencoders (VAEs ( Kingma et al. (2013))) map images to a Gaussian distribution in the latent space via an encoder, while a decoder reconstructs images from latent variables. Compared to traditional AEs, VAEs learn the data distribution rather than specific samples, mitigating overfitting. However, the generated images often appear blurry due to insufficient latent space constraints, as the KL regularization weight is relatively low. Vector Quantized Variational Autoencoders (VQ-VAEs ( Van Den Oord et al. (2017))) enhance compression efficiency by introducing a discrete latent space through vector quantization. The integration of compression techniques with diffusion models has given rise to various latent space-based diffusion models, with Latent Diffusion Models (LDMs) being the most prominent. Benefiting from the efficiency advantages of the latent space and the cross-modal cross-attention mechanism of LDMs, a diverse range of image generation(eg. Stable Diffusion( Rombach et al. (2022))) and image editing models(eg. DiffEdit( Couairon et al. (2022))) has emerged.
Beyond traditional image synthesis and editing, various other diffusion model-based tasks achieve computational efficiency through the introduction of latent spaces, too. For videos with high-dimensionality, complex temporal dynamics and large spatial variations, learning video distributions within a low-dimensional latent space has proven to be an effective method for efficiently generating high-quality videos. For instance, Video LDM (Blattmann et al., 2023), MagicVideo (Zhou et al., 2022), PVDM (Yu et al., 2023) and LVDM (He et al., 2022) apply the latent diffusion model paradigm to video generation. Video LDM focuses on high-resolution video generation. Initially, the LDM is pretrained exclusively on images; subsequently, turn the image generator into a video generator by introducing a temporal dimension to the latent space diffusion model and fine-tuning on encoded image sequences, i.e., videos. LVDM proposes hierarchical diffusion in the latent space such that longer videos with more than one thousand frames can be produced. To further overcome the performance degradation issue for long video generation, LVDM introduces conditional latent perturbation and unconditional guidance. You et al. (You et al., 2024) perform 3D graph diffusion in a low-dimensional latent space, which is learned through cascaded 2D–3D graph autoencoders for low-error reconstruction and symmetry-group invariance, resulting in training that is an order of magnitude faster. Motion Latent-based Diffusion (Chen et al., 2023c) is able to produce vivid motion sequences conforming to the given conditional inputs and substantially reduce the computational overhead in both the training and inference stages by performing diffusion process on the motion latent space. AudioLDM (Liu et al., 2023a), a text-to-audio system built on a latent space that learns continuous audio representations from contrastive language-audio pretraining latents, offers advantages in both generation quality and computational efficiency. L3DG (Roessle et al., 2024) utilizes a compressed latent space learned by a vector-quantized variational autoencoder, coupled with a sparse convolutional architecture to efficiently operate on room-scale scenes. As a result, the complexity of the costly 3D Gaussians generation process through diffusion is significantly reduced. TabSyn (Zhang et al., 2023a) attempt to utilize a diffusion model with a carefully crafted latent space in tabular data synthesis tasks, discovering that generation quality significantly improves while synthesis speed also increases. In reinforcement learning, condensing the archive into a single model while retaining the performance and coverage of the original collection of policies has proved challenging. Hegde et al. (Hegde et al., 2023) propose using latent diffusion models to distill the archive into a single generative model over policy parameters, achieving a compression ratio of 13x. RNAdiffusion (Huang et al., 2024) compresses token-level, biologically meaningful representations of RNA sequences into a set of fixed-length latent vectors and reconstructs RNA sequences from these latent variables. It utilizes a latent diffusion model to achieve controllable and efficiently translated RNA sequence generation.
3.1.2 Loss Formulation
In this section, we examine methods that enhance the efficiency of different loss formulations in diffusion models. For score matching (Hyvärinen & Dayan, 2005), we present approaches to reduce its computational costs. For rectified flow (Liu et al., 2022b), we explore how its carefully designed formulation enables straight-line sampling trajectories and subsequent improvements that further increase efficiency. We also include Flow Matching’s Optimal Transport approach (Lipman et al., 2022) in the rectified flow discussion, as it similarly achieves direct trajectory learning through linear parameterization of probability paths.
@@ 3.1.2 Loss Formulation
Score Matching. Compared to DDPM’s straightforward optimization in Eq.(6), although score matching Eq.(9) avoids the computation of the partition function $\bm{Z}_{\bm{\theta}}$, it still faces computational challenges, particularly in high-dimensional data. The computation of the trace of the Hessian matrix substantially increases the complexity as the dimensionality grows. Specifically, computing the trace requires many more backward passes than the gradient, making score matching computationally expensive for high-dimensional data.
@@
Therefore, to address the computational inefficient issue of training process, Song et al. (2020b) observed that one-dimensional problems are typically much easier to solve than high-dimensional ones. Inspired by the idea of the Sliced Wasserstein Distance (Rabin et al., 2012), they proposed Sliced Score Matching. The core idea of sliced score matching is to project both the score function of the model $\bm{s}_{m}(\mathbf{x};\bm{\theta})$ and the data $\bm{s}_{d}({\mathbf{x}})$ onto a random direction $\mathbf{v}$, and compare the differences along that direction. The sliced score matching objective is defined as:
@@
While effective for dimensionality reduction, score estimation still faces challenges in low data density regions where data samples are sparse.
Figure 4: Illustration of the rectified flow.
@@
Building upon sliced score matching, to address the issue of inaccurate score estimation in low data density regions, Song & Ermon (2019) introduces a novel generative framework that employs Langevin dynamics to produce samples based on estimated gradients of the data distribution $p_{\text{data}}(\mathbf{x})$. They proposed Noise Conditional Score Networks (NCSN) $s_{\theta}(\mathbf{x},\sigma)$, which jointly estimate scores across multiple noise-perturbed data distributions. By conditioning on a geometric sequence of noise levels $\sigma_{3}>\sigma_{2}>\sigma_{1}$, a single network learns to estimate scores for distributions ranging from highly smoothed $p_{\sigma_{3}}(\mathbf{x})$ that fill low-density regions to concentrated $p_{\sigma_{1}}(\mathbf{x})$ that preserve the structure of the original data manifold. This unified training approach enables robust score estimation across the entire data space. Following a similar derivation, as Song et al. (2021), Dockhorn et al. (2021) introduces Coupled Langevin Dynamics (CLD), redefining the score matching objective within the CLD framework. Unlike traditional score matching methods that inject noise directly into the data space, CLD simplifies the task by only requiring the model to learn the score of the conditional distribution $p_{t}(v_{t}\mid x_{t})$, where noise is injected into an auxiliary variable $v_{t}$ coupled with the data.
Figure 5: Illustration of data-dependent adaptive priors for diffusion processes across different modalities, demonstrating how tailored priors improve generation on quality.
Rectified Flow. As illustrated in Figure 4, Rectified Flow, proposed by (Liu et al., 2022b; Liu, 2022), introduces a method for training ordinary differential equation (ODE) models by learning straight transport paths between two distributions, $\pi_{0}$ and $\pi_{1}$.
The key idea is to minimize the transport cost by ensuring that the learned trajectory between these two distributions follows the most direct route, which can be computationally efficient to simulate. Unlike traditional diffusion models, which may follow roundabout paths, Rectified Flow leverages a simpler optimization problem to create a straight flow, thereby improving both training efficiency and the quality of the generated samples. Flow Matching (Lipman et al., 2022), as discussed in  §2.1.4, introduces the Optimal Transport (OT) approach as a significant advancement in generative modeling. The key innovation of the OT method lies in how it parameterizes conditional probability paths. It represents conditional probability paths in Eq(19) using Gaussian distributions with means $\mu_{t}(x)=tx_{1}$ and standard deviations $\sigma_{t}(x)=1-(1-\sigma_{\min})t$ that change linearly with time. This linear parameterization produces a vector field with constant direction, enabling transitions from noise to data along straight-line trajectories as well.
Building upon the foundation of Rectified Flow, InstaFlow (Liu et al., 2023c) applies the Rectified Flow concept to text-to-image generation, achieving a significant breakthrough. InstaFlow represents a major advancement in efficient diffusion models, which are capable of high-quality image generation in just one step. It applied Rectified Flow to large-scale datasets and complex models like Stable Diffusion, introduced a novel text-conditioned pipeline for one-step image generation, and achieved an FID score of 23.3 on MS COCO 2017-5k.
InstaFlow’s success highlights the potential of Rectified Flow in dramatically reducing the computational cost of diffusion models while maintaining high output quality.
3.1.3 Training Tricks
Training tricks encompass a range of practical strategies aimed at improving the efficiency, convergence, and sample quality of diffusion models by optimizing various aspects of the learning process. In this section, we explore three key approaches: data-dependent adaptive priors, which tailor initial distributions to specific data characteristics for faster convergence; extra loss supervision, which introduces additional loss terms to better align the model with data distributions and enhance generation quality; and noise schedule design, which governs the addition and removal of noise to streamline the diffusion process and accelerate sampling.